When three pumps, A, B, C, are running together, they can pump 3200 gal per hour. When only A and B are running, 1800 gal per hour can be pumped, when only A and C are running 2000 gal_per hour can be pumped. What is the pumping capacity of each pump?
_______________________________________________________
$6300 is invested, part of it at !0% and part of it at 6%. For a certain year, the total yield is $502.00 How much was invested at each rate?
How much was invested at 10%
How much was invested at 6%

Answer by mananth(16946)

A= xgal/h, B=ygal/h, C= zgal/h
x+y+z=3200....1
x+y=1800......2
x+z= 2000......3
multiply (2) by -1 and add to (1)
-x-y=-1800
add it to (1)
z=1400
In eq(3)
x+z=2000
plug z
x+1400=2000
x=600
...
Plug x & z in (1)
600+y+1400=3200
y=1200
.....................
Part I 10.00% per annum
Part II 6.00% per annum
x+y=6300 ------------------------1
10.00%x+ 6.00%y=$502.00
Multiply by 100
10x+6y=$50,200.00 --------2

Multiply (1) by -10
we get

-10x-10y=-63000.00

Add this to (2)

-4y=-$12,800.00
divide by -4
y=$3,200.00 investment at 6.00%
Balance $3,100.00 investment at 10.00%
m.ananth@hotmail.ca